OpenStudy (anonymous):

can someone please help integrate this? >< V=2π⋅∫02ex−e−xdx

OpenStudy (anonymous):



OpenStudy (anonymous):

did in the other question u posted adi ......

OpenStudy (anonymous):

-4pi + 2/e^2pi + 2e^2pi

OpenStudy (anonymous):

i dont get how you got that gahh sorry ><

sam (.sam.):

\[2\pi \int\limits_{0}^{2}e ^{x}-e ^{-x} dx\] =\[2\pi(\frac{1}{e ^{2}}+e ^{2}-2)\]

OpenStudy (anonymous):

so as of now u just plugged in for you integral of the 2 right? wherever x is?

sam (.sam.):

Substitute 2 and 0 to x into equation, then minus them